How do I find the two arithmetic means between -13 and 8?
Answer by dfrazzetto(277)

There is only one arithmetic mean:
add the numbers and divide by 2 (because in this case there are 2 numbers):
mean(n1,n2) = (n1+n2)/2
m = -13 + 8 / 2 = -5/2 = -2.5